A 0.415 g sample of KHP will be titrated with a NaOH solution with an assumed concentration of 0.100 M. 

Calculate the volume of NaOH that should neutralize the KHP sample.


1
Expert's answer
2022-04-27T04:42:02-0400

n(KHP)=0.415/204.22=2.03_10^-3 mol; C(NaOH)V(NaOH)=n(KHP), thus V(NaOH)= (2.03_10^-3)/0.100=0.0203 liters = 20.3 ml